OpenStudy (anonymous):

Determine which line is parallel to 6x + 6y = 6 1. 2y − 3 = 6x 2. 6y − 3 = x 3. 6x = y + 3 4. None of them. 5. 3x − 6 = −3y

OpenStudy (anonymous):

6x + 6y = 6 divide throughout by 6 x + y = 1 gives y = -x + 1 5. 3x − 6 = −3y 3y = -3x + 6 divide throughout by 3 y = -x + 2 so 5. 3x - 6 = -3y will be parallel to 6x + 6y = 6 as both have same slope